OpenStudy (anonymous):

How do i find or approximate the points that are guaranteed to exist by the mean value theorem on f(x)=ln2x; [1,e] ?

OpenStudy (anonymous):

\[f'(c)=\frac{f(b)-f(a)}{b-a}\] \[a=1, b=e, f(1)=\ln(2), f(e)=\ln(2e)=\ln(2)+\ln(e)=\ln(2)+1\] \[f'(x)=\frac{1}{x}\]

OpenStudy (anonymous):

so your job is to solve for c \[\frac{1}{c}=\frac{\ln(2)+1-\ln(2)}{e-1}\] \[\frac{1}{c}=\frac{1}{e-1}\]

OpenStudy (anonymous):

alright, i'll get on that

OpenStudy (anonymous):

one step left! take the reciprocal and you are done

OpenStudy (anonymous):

okay, i'll piece it together from here, thanks

OpenStudy (anonymous):

i just gotta look it all over

OpenStudy (anonymous):

the statement of the mvt is easy to read, the hard part is in going from the general \[f'(c)=\frac{f(b)-f(a)}{b-a}\] to the specific

OpenStudy (anonymous):

so it would just be c=e-1?

OpenStudy (anonymous):

or c=(c^2)/(e-1)

OpenStudy (anonymous):

i got \[c=e-1\]